Find each probability if a die is rolled 4 times
P(exacty one 3)
P(exacty three 3s)
P(at most two 3s)
P(at least 3s)

Answer by edjones(8007)

Let x=1/6 and y=5/6
(x+y)^4=x^4+4x^3y+6x^2y^2+4xy^3+y^4
P(exacty one 3)
4xy^3=.3858
.
P(exacty three 3s)
4x^3y=.0154
.
P(at most two 3s)
y^4+4xy^3+6x^2y^2=.9838
.
Ed